What is the answer to the expression
(8/27) to the (-4/3) power

Answer by stanbon(75887)

What is the answer to the expression
(8/27) to the (-4/3) power
-----
Look at the exponent.
The 3 in the denominator means "take the cube root" to get:
= (2/3)^-4
-------
The negative means "invert the fraction" to get:
= (3/2)^4
-------
Now raise that to the 4th power to get:
= 81/16
==============
Cheers,
Stan H.
================